A farmer wants to enclose a rectangular lot of using 48 meters of fencing materials. If the wall acts as one side of the fence, what is the maximum area that can be fenced?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

length = L
width = W

Assuming the wall is one length, then we have 2W + L = 48. This is maximized when W = L.

x = side length
3x = 48
x = 16

area = x^2 = 256 m^2

Answer by ikleyn(52781)

.


            The answer and the solution by @ThePenguin are both incorrect.



The correct answer is 12*24 = 288 square meters for the area.

The optimum dimensions are 12 m width and 24 meters length (along the wall).

See the lesson
    - A farmer planning to fence a rectangular area along the river to enclose the maximal area
in this site.